# Orbital Depot Location Optimization for Satellite Constellation Servicing with Low-Thrust Transfers

### v1

@@ 3.2 Nonlinear Programming with Fixed Routing Solution
3:for $n=1$ to $n_{\max}$ do
4:     Solve the routing subproblem $P_{L}$ to obtain an updated routing solution ${\bf\it z}_{\text{new}}$
5:            using the current depot locations ${\bf\it\eta}$ and the warm start ${\bf\it z}$
6:     Solve the depot location subproblem $P_{NL}$ to obtain new depot locations ${\bf\it\eta}_{\text{new}}$
7:            using the updated routing solution ${\bf\it z}_{\text{new}}$
8:     if $\|{\bf\it\eta}_{\text{new}}-{\bf\it\eta}\|<\texttt{tolerance}$ then
9:         Return depot locations ${\bf\it\eta}_{\text{new}}$ and routing solution ${\bf\it z}_{\text{new}}$
10:     end if
11:     Update depot locations: ${\bf\it\eta}\leftarrow{\bf\it\eta}_{\text{new}}$
12:     Update warm start solution: ${\bf\it z}\leftarrow{\bf\it z}_{\text{new}}$
13:end for
14:Return final depot locations ${\bf\it\eta}$, routing solution ${\bf\it z}$, and a message: Max iterations reached
@@ 4 Case Study
This section demonstrates the mathematical formulation and solution framework of the CLRPOD through a case study focusing on a servicing scenario for a satellite constellation. For simplicity, all satellites’ and depots’ orbits are approximated as circular orbits so that only three orbital elements (i.e., semimajor axis, inclination, and RAAN) are sufficient to determine each orbit. Each orbital depot is assumed to store necessary fuel and equipment, and a servicer corresponding to each depot departs from the depot, visits multiple satellites, performs in-space servicing, and returns to the depot. If it is necessary, the servicer will depart again for other satellites.
@@
Virtual orbit radius for EMLEO calculation, $r_{0}$, km 6,578
Minimum orbit radius available for the depot, $r_{\text{min}}$, km 6,578
@@ 4.1 Test Case with GPS Constellation
This case study focuses on the servicing of the GPS satellites, consisting of 18 satellites ($n_{t}=18$) and three depots ($n_{d}=3$). The GPS satellites are positioned in MEO at a radius of approximately 26,560 km. They have nearly circular orbits, with similar inclinations of around $55^{\circ}$, and are separated by $60^{\circ}$ in RAAN. Detailed orbital elements are provided in Table A1. Note that eccentricity and the argument of perigee are not included, as we assume the orbits are circular. The initial locations of the depots are manually set to be approximately middle values of the satellites’ orbital elements, which is intuitively reasonable.
Figures 2 and 3 illustrate the comparisons between the solution using the initial depot locations and the solution derived from the CLRPOD framework. In these figures, squares represent depots, triangles denote the GPS satellites, black arrows indicate the visiting sequence for servicing, and the green circle in Fig. 2 marks the minimum radius allowed for the depots.
@@
Table 3 summarizes the case study results. Because the specific impulses of the launch vehicle and depot are significantly lower than that of the servicer, the depots are located at the lowest possible altitudes to minimize overall propellant consumption. Additionally, the depots are positioned with inclinations of approximately $50^{\circ}$ instead of $55^{\circ}$. This adjustment is necessary because the angle difference between orbital planes ($\Delta\theta$) is a function of both inclination and RAAN. By fine-tuning both the inclinations and RAAN, we can reduce the plane-change angle, directly affecting the amount of propellant consumed. As a result, the total propellant required in EMLEO is reduced from 8,255.936 kg to 5,197.532 kg, yielding a 37.04% reduction compared to the initial guess. This reduction from the initial guess quantitatively demonstrates the effectiveness of optimizing the depot locations. The solution framework converges in a single iteration, resulting in a computation time of 203.341 seconds.
@@
Figure 2: Comparison of the ISAM solutions for this case study ($a$-$\Omega$ plot): a) Solution with initial depot location guesses; b) Solution with final depot locations from the proposed framework.
@@
Figure 3: Comparison of the ISAM solutions for this case study ($i$-$\Omega$ plot): a) Solution with initial depot location guesses; b) Solution with final depot locations from the proposed framework.
@@
Depot Index Initial Guess Final Solution Computation Time, s
@@
1 26,560.0 55.00 0.00 6578.0 52.63 351.69 203.341
2 26,560.0 55.00 120.00 6578.0 49.86 110.91
3 26,560.0 55.00 240.00 6578.0 50.22 232.71
Total Propellant EMLEO, kg 8,255.936 5,197.532 (37.04 % $\downarrow$)
@@ 4.2 Numerical Experiments
Comprehensive numerical experiments are conducted to evaluate the performance of the CLRPOD and its solution procedure, including the EMLEO reduction from the initial guess, the number of interactions and the computational time. In this study, 100 problem instances are generated for each of five different numbers of depots (with $n_{d}=1,2,3,4,$ and $5$) and four different numbers of satellites (with $n_{t}=5,10,15,$ and $20$). This results in a total of 2,000 problems being solved. The semimajor axes of all satellites are randomly generated in the range from 7,000 km (LEO) to 42,164 km (GEO). Both the inclination and the RAAN are varied within the ranges of $[0^{\circ},180^{\circ}]$ and $[0^{\circ},360^{\circ}]$, respectively. The initial guess for the depot locations is automatically generated using k-means clustering, which gives fast and intuitive results as we already know how many depots are needed. However, as demonstrated in the earlier GPS case study, the proposed solution framework can also handle any initial guesses obtained from other methods.
@@
Number of iterations, - Computation time, s Solve success rate, %
@@ 5 Conclusions
Appendix
@@ Acknowledgments
This work was conducted with support from the Air Force Office of Scientific Research (AFOSR), as part of the Space University Research Initiative (SURI), under award number FA9550-23-1-0723. I thank Yuri Shimane for reviewing and verifying my code. AI technologies were used for grammar checking and refinement.

### v2

@@ 3.2 Nonlinear Programming with Fixed Routing Solution
3:for $\texttt{iter}=1$ to $\texttt{iter}_{\max}$ do
4:  Solve the routing subproblem $P_{L}$ to obtain an updated routing solution ${\bf\it z}_{\text{new}}$ using the current depot locations $\bf\it\eta$} and the warm start $\bf\it z$}
5:  Solve the depot location subproblem $P_{NL}$ to obtain new depot locations ${\bf\it\eta}_{\text{new}}$ using the updated routing solution ${\bf\it z}_{\text{new}}$
6:  if $\|{\bf\it\eta}_{\text{new}}-{\bf\it\eta}\|<\texttt{tolerance}$ then
7:   Return depot locations ${\bf\it\eta}_{\text{new}}$ and routing solution ${\bf\it z}_{\text{new}}$
8:  end if
9:  Update depot locations: ${\bf\it\eta}\leftarrow{\bf\it\eta}_{\text{new}}$
10:  Update warm start solution: ${\bf\it z}\leftarrow{\bf\it z}_{\text{new}}$
11:end for
12:Return final depot locations $\bf\it\eta$}, routing solution $\bf\it z$}, and a message: Max iterations reached
@@ 4 Case Study
This section demonstrates the mathematical formulation and solution framework of the CLRPOD through a case study focusing on a servicing scenario for a satellite constellation. For simplicity, all satellites’ and depots’ orbits are approximated as circular orbits so that only three orbital elements (i.e., semimajor axis, inclination, and RAAN) are sufficient to determine each orbit. Since we assume a static problem with regular satellite servicing (i.e., no emergency or time-critical tasks), all services can be fully planned in advance. Accordingly, each depot carries the necessary amount of fuel to its optimized location, as determined by the proposed solution framework. A servicer corresponding to each depot departs from the depot, visits multiple satellites, performs in-space servicing, and returns to the depot. If it is necessary, the servicer will depart again for other satellites.
@@
Virtual orbit radius for EMLEO calculation, $r_{0}$, km 7,000
Minimum orbit radius available for the depot, $r_{\text{min}}$, km 7,000
@@ 4.1 Test Case with GPS Constellation
This case study focuses on the servicing of the GPS satellites, consisting of 18 satellites ($n_{t}=18$) and three depots ($n_{d}=3$). The GPS satellites are positioned in MEO at a radius of approximately 26,560 km. They have nearly circular orbits, with similar inclinations of around $55^{\circ}$, and are separated by $60^{\circ}$ in RAAN. Detailed orbital elements are provided in Appendix A (Table A1). Note that eccentricity and the argument of perigee are not included, as we assume the orbits are circular. Additionally, the actual RAAN of depots may vary due to J2 perturbation; see Appendix B for the corresponding sensitivity analysis. The initial depot locations are determined using k-means clustering, where the centroids of the resulting clusters serve as the initial guesses for the depot positions.
Figures 2 and 3 illustrate the comparisons between the solution using the initial depot locations and the solution derived from the CLRPOD framework. In these figures, squares represent depots, triangles denote the GPS satellites, black arrows indicate the visiting sequence for servicing, and the green circle in Fig. 2 marks the minimum radius allowed for the depots. Note that the initial depot locations obtained from k-means clustering may not appear well-centered with respect to the satellites they ultimately serve. This is due to the fundamental difference between k-means clustering and routing optimization objectives. While k-means minimizes intra-cluster variance, the routing optimization seeks to minimize overall propellant consumption. As a result, satellites may be serviced by depots outside their original k-means clusters. Nevertheless, in certain configurations, as seen in the numerical results, the optimized depot locations can coincide with the initial k-means centroids, resulting in negligible or zero EMLEO reductions.
@@
Table 3 summarizes the case study results. Because the specific impulses of the launch vehicle and depot are significantly lower than that of the servicer, the depots are located at the lowest possible altitudes to minimize overall propellant consumption. However, the placement of depots at the lowest altitude is not universally optimal. Higher-altitude depot locations may yield better overall efficiency depending on mission-specific factors such as satellite distribution, servicing sequence, and propulsion characteristics. Additionally, the depots are positioned with inclinations of approximately $50^{\circ}$ instead of $55^{\circ}$. This adjustment is necessary because the angle difference between orbital planes ($\Delta\theta$) is a function of both inclination and RAAN. By fine-tuning both the inclinations and RAAN, we can reduce the plane-change angle, directly affecting the amount of propellant consumed. As a result, the total propellant required in EMLEO is reduced from 7,773.982 kg to 4,906.056 kg, yielding a 36.891% reduction compared to the initial guess. This reduction from the initial guess quantitatively demonstrates the effectiveness of optimizing the depot locations. The solution framework converges in two iterations, resulting in a computation time of 305.923 seconds.
@@
Figure 2: Comparison of the in-space servicing solutions for this case study ($a$-$\Omega$ plot): a) Solution with initial depot location guesses; b) Solution with final depot locations from the proposed framework.
@@
Figure 3: Comparison of the in-space servicing solutions for this case study (ii-$\Omega$ plot): a) Solution with initial depot location guesses; b) Solution with final depot locations from the proposed framework.
@@
Depot Index Initial Guess Final Solution
Computation
Time, s
@@
1 26,560.32 55.65 317.28 7,000.00 51.59 296.41 305.923
2 26,572.91 55.39 17.68 7,000.00 51.87 33.04
3 26,560.14 54.51 151.08 7,000.00 50.92 171.41
Total Propellant EMLEO, kg 7,773.982 4,906.056 (36.891 % $\downarrow$)
@@ 4.2 Numerical Experiments
Comprehensive numerical experiments are conducted to evaluate the performance of the CLRPOD and its solution procedure, including the EMLEO reduction from the initial guess, the number of interactions and the computational time. In this study, 100 problem instances are generated for each of five different numbers of depots (with $n_{d}=1,2,3,4,$ and $5$) and four different numbers of satellites (with $n_{t}=5,10,15,$ and $20$). This results in a total of 2,000 problems being solved. The semimajor axes of all satellites are randomly generated in the range from 7,000 km (LEO) to 42,164 km (GEO). Both the inclination and the RAAN are varied within the ranges of $0^{\circ},180^{\circ}]$ and $0^{\circ},360^{\circ}]$, respectively. The initial guess for the depot locations is automatically generated using k-means clustering, which gives fast and intuitive results as we already know how many depots are needed. However, as demonstrated in the earlier GPS case study, the proposed solution framework can also handle any initial guesses obtained from other methods.
@@
Number of iterations, - Computation time, s
Solve success rate, %
@@ 5 Conclusions
Appendix A: Data Table Used in Case Study
@@
Appendix B: Sensitivity Analysis on RAAN Drift due to J2 Perturbation
In this study, all orbits are assumed to be fixed, without considering J2 perturbation. However, in reality, the J2 effect causes the changes in the RAAN depending on orbital altitude. For example, the RAAN of a GPS orbit changes by approximately $0.0387^{\circ}/\text{day}$, whereas an orbit with a radius of 7,000 km experiences a drift of approximately $4.121^{\circ}/\text{day}$.
Given the use of a low-thrust propulsion system, each orbital transfer is assumed to take approximately 10–50 days, and the overall mission (servicing 3 to 4 satellites per route) would span roughly 100 days. Therefore, the impact of J2 perturbation is not negligible, and a depot’s RAAN may drift significantly throughout the mission.
To evaluate the sensitivity of the optimization results to RAAN drift, we conducted a sensitivity analysis by varying the RAAN of the depot orbit. Figure B1 presents the variation in the objective value (total propellant consumption in terms of EMLEO) as a function of the depot’s RAAN for the GPS case study. It is important to note that we evaluate relative RAAN differences between depots and GPS satellites; thus, only the return transfer to the depot is affected by RAAN drift. The red dot in each figure represents the initial RAAN value (i.e., the case in which no RAAN drift occurs and the depot remains fixed).
The results show that the objective value changes by an average of 6.62%, with a maximum deviation of 14.98% from the original solution. This level of variation is considered acceptable, as the proposed method is intended for high-level strategic planning rather than detailed tactical trajectory optimization.
It is also important to note that solutions generated by the k-means clustering would be similarly affected by RAAN drift. Therefore, the observed variation does not imply that our proposed method is more sensitive to J2 effects than the k-means clustering. The performance improvements over the k-means clustering remain valid.
Figure B1: EMLEO variation due to RAAN drift of: (a) Depot 1; (b) Depot 2; (c) Depot 3.
@@ Acknowledgments
This work was conducted with support from the Air Force Office of Scientific Research (AFOSR), as part of the Space University Research Initiative (SURI), under award number FA9550-23-1-0723. We thank Yuri Shimane for reviewing and verifying the code. AI technologies were used for grammar checking and refinement.